ID |
Date |
Author |
Type |
Category |
Subject |
12073
|
Wed Apr 13 00:56:07 2016 |
gautam | Update | endtable upgrade | X endtable repopulation |
Over the last couple of days, I've been working on restoring the optical layout on the X-endtable. Some notes about the status as of today:
Lightwave NPRO output power
The output power from the lightwave NPRO is about 210mW (as measured with the calorimeter). This is significantly lower than the value of ~300mW reported in this elog. It may be that the laser crystal temperature has changed compared to that measurement, but the "ADJ" parameter is at 0, both today and in that measurement. The laser has also been on for more than a day now, that should be sufficient time for the crystal to equilibriate to its final operating state? Is such a large change in output power possible just because of a change in laser crystal temperature? Or did the laser really lose ~1/3rd of its output power over the last two months?
Alignment into IR Faraday, and changes to the planned layout
I've set up the layout until steering the beam through the IR faraday. The input power into the IR Faraday is ~210mW. The output power is ~186mW, after optimizing the angle of the HWP. These numbers seem consistent with what I had reported in this elog (although this was for the Innolight NPRO). The alignment looks reasonably good to the eye as well.
I've made one change to the planned layout (latest version here). Y1 is now a 2" 99% reflective for S polarization beam splitter, instead of a 1" HR mirror. I made this change because we want some light from the NPRO to be transmitted through this optic to couple into the fiber eventually, for the IR beat. I measured the transmitted power to be ~1.5mW, which is around what we were coupling into the fiber before, and should suffice now. The Lightwave NPRO datasheet (page 4) suggests that the polarization of the output of the laser is S, and the measured power before and after this optic suggests that it is working as advertised. This means that HWP 1 also has to be moved downstream (to rotate the polarization so as to maximize transmission through the IR faraday). Space constraints meant that I could not mount HWP 1 on the baseplate+3/4" OD post assembly which is what we want where possible on the new table, so for this optic, I used a 1" OD post and a fork. There may be a couple of other optics in the final layout where space constraints dictate we compromise in this way.
I've also installed beam dumps for the rejected light from the Faraday. For now, these are the old beam dumps. They looked reasonably intact. I believe we have a bunch of new beam dumps on hand as well, so these can be swapped out if deemed necessary.
Cleaning of optics
All the optics are being cleaned using first contact before being installed on the table.
As I found out the hard way, it is not a good idea to clean small optics like half-wave plates while in their mounts. The first contact tends to bond to the frame while drying, and doesn't come off cleanly. Koji helped me clean the offending pieces (he used tweezers to manually remove the residual first contact, and then some acetone to clean up any remaining residue). Subsequently, he re-cleaned these optics, again using first contact, but this time being careful not to extend all the way out to the edge of the optic. The idea is to cover as much area as possible with first contact, while staying clear of the edge. This approach worked reasonably well.
The next major step is to achieve optimal alignment into the doubler. I've placed the doubler on the table in it's approximate final position, I wanted to make sure the enclosure support wasn't in the way (it isn't). The cable from the oven won't run all the way to the Thorlabs temperature controller in it's usual place, we need to either extend the cable, or figure out a new place where we can keep the temperature controller. |
12072
|
Tue Apr 12 22:41:00 2016 |
Koji | Update | IOO | PMC/IMC aligned, WFS offset adjusted |
Did it again.
PMC Trans ~0.739
IMC Trans ~15000 |
12071
|
Tue Apr 12 09:14:57 2016 |
Steve | Update | SUS | Ruby wire - v - groove cut pictures |
The ruby wire standoff V groove cuts are looking good.
I will request free sample of sapphire prizm where one side would have SOS's R cylindrical surface.
The present plan to have the v-groove on this prism.
|
Attachment 1: Sapphire_prism_wire_standoff.JPG
|
|
12070
|
Mon Apr 11 17:03:41 2016 |
Steve | Update | Calibration-Repair | 1W Innolight repair completed |
The laser is back. Test report is in the 40m wiki as New Pump Diode Mephisto 1000
It will go on the PSL table. |
12069
|
Mon Apr 11 16:06:30 2016 |
ericq | Update | LSC | DRFPMI Data Archived |
I have copied over the complete frame files from two DRFPMI lock acquisitions + locks to /frames/archive . The data should be safe from the wiper script here.
One, under the subfolder DRFPMI_Mar29_cal is the lock where the CAL-DARM channel is properly calibrated at GPS time 1143274087.
The other lock, under DRFPMI_MAR29_nocal , does not have the calibration set up yet, but was a much quicker acquistion (<2 min from ALS acquisition to DRFPMI) and longer lock (~8min). |
12068
|
Fri Apr 8 09:13:20 2016 |
Steve | Update | endtable upgrade | ETMX-T beam height |
X arm resonating after alignment, beam height on ETMX optical table ~4.75"
|
Attachment 1: ETMXalignedT.png
|
|
Attachment 2: ETMXF.jpg
|
|
12067
|
Thu Apr 7 14:30:58 2016 |
Steve | Update | endtable upgrade | ETMX table height |
Quote: |
Steve has finished installing the enclosure on the new endtable. So Eric and I decided to try and lock the X arm and measure the beam height of the transmitted IR beam relative to the endtable. We initially thought of using POX DC as a the LSC trigger but this did not work as there was no significant change in it when the arm was flashing. Eric then tried misaligning the ITM and using AS110 as a trigger - this worked. We then recompiled the ASS model to take AS110 as an input, and ran the dither alignment. After doing so, I measured the beam height at two points on the new endtable.
Bottom line:
- The beam is roughly level across the table (along the North-South direction, within the precision to which I could place the irides and measure the height). The table has also been levelled pretty well...
- The beam height is ~4.7" across the endtable
So the beam is about 0.7" higher relative to the endtable than we'd like it to be. What do we do about this?
- Is it even possible to raise the table by 0.7" so we can have a level beam everywhere? Are there some constraints related to how the enclosure is attached to the window?
- Are we okay with tolerating a solution where we keep the beam level at 4", and use Y10 and Y11 (see layout in elog 12060) to raise the beam by 0.7", and then have slightly higher posts for the optics downstream of this point?
I've also placed two irides extending the cavity axis on the endtable. These should be helpful in aligning the green to the arm eventually.
|
The new TMC 4' x 3' x4" optical table and enclosure is installed - aligned- leveled.
Atm2, Picture is taken ~42" from the window at 3.75 camera height. The leveled table height is wthin 1/4 at the center of the window.
I think this is close enough to move on with the installation of the optics.
We can raise the loaded table in the future if it is needed.
Atm4, Optical table height to floor 33" at the south west corner
Atm3, Enclosure top cover transmission at 1064 nm, 1mm beam size, power level 157 mW, 0 degree incident angle, T 1.3% Metal shield is required above 100 mW hitting the wall of the enclosure!
Atm5, window to enclosure Kapton seal |
Attachment 1: ETMXbrs.png
|
|
Attachment 2: ETMX.jpg
|
|
Attachment 3: IMG_0618.JPG
|
|
Attachment 4: ETMX.jpg
|
|
Attachment 5: ETMXwc.jpg
|
|
12066
|
Thu Apr 7 12:51:24 2016 |
gautam | Update | endtable upgrade | Beam height differences |
Steve has finished installing the enclosure on the new endtable. So Eric and I decided to try and lock the X arm and measure the beam height of the transmitted IR beam relative to the endtable. We initially thought of using POX DC as a the LSC trigger but this did not work as there was no significant change in it when the arm was flashing. Eric then tried misaligning the ITM and using AS110 as a trigger - this worked. We then recompiled the ASS model to take AS110 as an input, and ran the dither alignment. After doing so, I measured the beam height at two points on the new endtable.
Bottom line:
- The beam is roughly level across the table (along the North-South direction, within the precision to which I could place the irides and measure the height). The table has also been levelled pretty well...
- The beam height is ~4.7" across the endtable
So the beam is about 0.7" higher relative to the endtable than we'd like it to be. What do we do about this?
- Is it even possible to raise the table by 0.7" so we can have a level beam everywhere? Are there some constraints related to how the enclosure is attached to the window?
- Are we okay with tolerating a solution where we keep the beam level at 4", and use Y10 and Y11 (see layout in elog 12060) to raise the beam by 0.7", and then have slightly higher posts for the optics downstream of this point?
I've also placed two irides extending the cavity axis on the endtable. These should be helpful in aligning the green to the arm eventually. |
12065
|
Wed Apr 6 17:52:21 2016 |
gautam | Update | endtable upgrade | First contact cleaning commenced |
I've begun cleaning the optics that will eventually go back onto the newly installed X-endtable. We decided that First Contact was the way to go (as opposed to methanol drag wiping). Koji demonstrated the application of the (red) First Contact solution onto a 2" mirror - I then proceeded to work on the rest of the optics. We are broadly following the procedure in E1000079 - first one coat of First Contact solution is applied, then a small piece of PEEK is embedded by applying a second layer of solution over it (this will enable us to pull off the First Contact once we are ready - the plan is to do this after roughly placing the optic on the table. As of now, I've finished coating most of the optics that are part of the IR Transmon path - I will continue later in the evening.
The new endtable is almost ready for re-population. Steve just needs to shim the enclosure which will be done tomorrow morning. The game-plan as discussed at the meeting today is to first try and set up the IR Transmon path. This will allow us to verify that the endtable height is such that we can maintain a beam height of 4" everywhere on the table (I suspect we may have to compromise at some poing and do some fine adjustment of 1/4 to 1/2" somewhere though). It will also allow me to define the cavity axis relative to the table, which will be useful to place the green steering optics eventually. Doing this will be challenging though as right now, I can't see any of the arm flashes on the endtable using an IR card. Ideally, we want to somehow lock the X arm and then do the checks mentioned at the endtable, before beginning to put the endtable back together. |
12064
|
Tue Apr 5 14:16:34 2016 |
gautam | Update | CDS | BLRMS for optics suspensions - library block UPDATED |
As discussed in a Wednesday meeting some time ago, we don't need to be writing channels from BLRMS filter modules to frames at 16k (we suspect this is leading to the frequent daqd crashes which were seen the last time we tried setting BLRMS up for all the suspensions). EricQ pointed out to me that there conveniently exists a library block that is much better suited to our purposes, called BLRMS_2k. I've replaced all the BLRMS library blocks in the sus_single_BLRMS library block that I made with there BLRMS_2k blocks. I need to check that the filters used by the BLRMS_2k block (which reside in /opt/rtcds/userapps/release/cds/common/src/BLRMSFILTER.c) are appropriate, after which we can give setting up BLRMS for all the suspensions a second try... |
12063
|
Tue Apr 5 11:42:17 2016 |
gaericqutam | Update | endtable upgrade | TABLE REMOVAL |
There is currently no table at the X end!
We have moved the vast majority of the optics to a temporary storage breadbord, and moved the end table itself to the workbench at the end.
Steve says Transportation is coming at 1PM to put the new table in. |
12062
|
Tue Apr 5 08:55:51 2016 |
Steve | Update | SUS | local EQ 3.1m |
Local earth quake 3.1 magnitude in Valencia, Ca did not trip our suspensions.
|
Attachment 1: eq3.1Valencia.png
|
|
12061
|
Mon Apr 4 15:04:14 2016 |
gautam | Update | endtable upgrade | COMPONENT REMOVAL |
I'm planning to start removing components from the X endtable tomorrow morning at ~10AM - if anyone thinks I should hold off and do some further checks/planning, let me know before this so that I can do the needful.
|
12060
|
Mon Apr 4 10:59:12 2016 |
gautam | Update | endtable upgrade | proposed layout v3 |
I realized I had overlooked an important constraint in the layout, which is that the enclosure will have two supports that occupy some region of the table - these are denoted in blue in v3 of the layout (Attachment #1). I measured the dimensions for these from the existing Y-endtable. The main subsystem this has affected is the IR transmission monitors, but I've been able to move the photodiodes a little to accommodate this constraint.
I've also done the mode-matching calculations explicitly for the proposed new layout (Attachments #2 and #3, code in Attachment #4). While the layout was largely adopted from what Andres posted in this elog, I found that some of the parameters he used in his a la mode code were probably incorrect (e.g. distance between the 750mm lens and the ETM). More critically, I think the Gouy phase for the optimized solution in the same elog is more like 60 degrees. I found that I could get a (calculated) Gouy phase difference between the two PZT mirrors of ~81 degrees by changing the green path slightly, and making the two PZT mirrors Y7 and Y8 (instead of Y7 and Y11, for which the Gouy phase difference is more like 50 degrees). But this way the two steering mirrors are much closer to each other than they were before. Other misc. remarks about the mode matching calculations:
- The beam diameter at the locations where the Faraday isolators should go is well below 5mm, the aperture size of the Faraday isolators
- The calculated mode-matching efficiencies suggest that we don't need any cylindrical lenses though the mode from the NPRO is elliptical
- Attachment #5 is a CAD drawing of the layout with all dimensions used for the mode-matching calculations included (although they are in inches)
These changes also necessitated minor changes to the transmitted IR beampath and the Oplev system, but these changes are minor. I've also switched the positions of the AUX IR power monitoring PD and the fiber coupler as suggested by Koji. The shutter has also been included. |
Attachment 1: ETMX_proposed_layout_v3.pdf
|
|
Attachment 2: IR_modematch.pdf
|
|
Attachment 3: Green_modematch.pdf
|
|
Attachment 4: XendALaMode.zip
|
Attachment 5: ETMX_proposed_layout_dimensions.dwg
|
12059
|
Fri Apr 1 13:11:26 2016 |
ericq | Update | WienerFiltering | Bilinear Noise Testing |
I've been banging my head against bilinear noise subtraction, and figured I needed to test things on some real hardware to see if what I'm doing makes sense.
I ran the ASS dither alignment on the Y arm, which ensures that the beam spots are centered on both mirrors.
I then drove ITMY in yaw with some noise bandpassed from 30-40 Hz. It showed the expected bilinear upconversion that you expect from angular noise on a centered beam, which you can see from 60-80 Hz below
I looked at the length signal, as the noise subtraction target, and the ITMY oplev yaw signal plus the transmon QPD yaw signal as witnesses.

There is some linear coupling to length, which means the the centering isn't perfect, and the drive is maybe large enough to displace it off center. However, the important part is the upconverted noise which is present only in the length signal. The QPD and oplev signals show no increased noise from 60-80Hz above the reference traces where no drive is applied
I then compared the multicoherence of those two angular witnesses vs. the multicoherence of the two (linear) witnesses plus their (bilinear) product. Including the bilinear term clearly shows coherence, and thereby subtraction potential, at the upconverted noise hump.

So, it looks like the way I'm generating the bilinear signals and calculating coherence in my code isn't totally crazy. |
Attachment 1: bilinear_drive.pdf
|
|
Attachment 2: 40m_bilin.pdf
|
|
12058
|
Thu Mar 31 19:49:31 2016 |
gautam | Update | endtable upgrade | proposed layout v2 |
The major changes from the previous layout:
- I've depicted the Green reflected beam path more accurately - I approximately measured the angle of the rejected beam from the faraday from the Y-end setup. This looks like a workable solution, and is similar to what we have currently at the Y-end
- I've added some optics to monitor the DC power and RIN of the AUX laser
- I've added two lenses to the input path of the Oplev beam (the path is such that I think we can use the same lenses that are currently being used.
- I've now drawn the beams in CAD so that is marginally neater.
To do:
- Post mode matching solutions for AUX laser to doubler and green beam to arm for this proposed layout (should be identical to what we have now, which at least according to the calculation is a good solution, but I will double check - I also need to quantify what the effect of the elliptical beam is)
- Check the Gouy phase of the transmitted IR beam at the QPD - we may need to change some lenses in this path. But I think the path as such is close enough (distance-wise) to what we have currently at the X end (after accounting for the fact that the new endtable edge will be closer to the ETM) so I don't expect this to be a show-stopper.
Does any part of this layout need a radical redesign? |
Attachment 1: ETMX_proposed_layout_v2.pdf
|
|
12057
|
Thu Mar 31 09:38:41 2016 |
Steve | Update | endtable upgrade | Y end 4x3 existing layout |
Beam colors: 1064 nm red, 514 nm green and 633 nm yellow.
There should be room for lens in front of the pd at red3 and a mirror for alignment in the new layout.
This picture may help you how to improve the new ETMX 4' x 3' optical layout.
|
Attachment 1: ETMY4x3layout.jpg
|
|
12056
|
Wed Mar 30 17:38:52 2016 |
gautam | Update | endtable upgrade | X end table proposed layout |
Attachment 1: This is a photo of the current X end table optical layout with the beampaths of the various sub-systems overlaid. For the labels, see Attachment #2.
Attachment 2: This is a summary of all the optical components that are currently being used. I've noted some things we may want to change when we effect the swap. The important ones are:
- Switch out all 1" and 2" optic mounts which are not of the Polaris type to the Polaris type. I have checked that we have sufficient numbers of these in hand.
- Adjust the collimating lens of the fiber collimating telescope to get a better mode
- Many of the labels are probably outdated, now would be a good time to update them
- For the mode-matching of the AUX IR into the doubling crystal, a la mode suggests a better (i.e. less sensitive to lens position) solution is effected with L2 as a 100mm fl lens rather than 88.3mm. I did not change this during the laser swap in order to minimize the number of components changed. Since we are doing a wholesale change now, it may not be a bad idea to swap this out as well. I have checked that we have a suitable AR1064 coated lens.
- Some optics probably need to be cleaned...
- PZT mirror 2 has a new mount ready that is the "correct" height so we don't have to keep using makeshift stacked posts.
- The plan as it stands is to use the green coloured mount for the IR faraday (IO-5-1064-HP).
Have I missed anything important?
Attachment #3: I've made a CAD drawing of the proposed new layout and have overlaid the beampath in an amateur way because I couldn't figure OptoCad out - I figure this will suffice for now. I have adopted elements from the current Y-end layout, but have used Anders' mode-matching solution (same lenses, same positions of optics) to make sure we have good Guoy phase separation between the two PZT steering mirrors. Some notes:
- I've tried to palce the optics for the AUX IR into the doubler and subsequent steering of green into the arm cavity as per the mode matching solution. These should be pretty accurate, and the layout suggests we have some room to maneuver
- The Green REFL beampath is exaggerated but I think we have enough room to place Y16 appropriately and steer the reflected beam into the PDA36A
- We need two more 1" 1064nm coated mirrors for the initial steering into the doubling oven, I have checked we have these in hand.
- The IR pickoff into the fiber coupler may change somewhat once we change the mode and redo the mode-matching calculations. But again, I think we have sufficient room to implement a workable solution.
- After accounting for the fact that the new endtable will be a little closer to the vacuum chamber, Y12 in the proposed layout will be ~10cm further away from ETMX than it is currently. But as discussed at the meeting today, the Rayleigh range of the green beam should be large enough here such that this shouldn't be a significant change.
Steve says the table is ready - so if we are happy with this layout, we can move forward... |
Attachment 1: ETMX_3x2.JPG
|
|
Attachment 2: layout_details_20160328.pdf
|
|
Attachment 3: ETMX_proposed_layout.pdf
|
|
12055
|
Wed Mar 30 16:40:24 2016 |
ericq | Update | LSC | 2016 vs 2010 |
I haven't found any data files for the DARM spectrum of the previous generation of 40m, but with some GIMP-fu, I have plotted Monday's spectrum (green) on top of one of the figures from Rob's thesis.

|
12054
|
Wed Mar 30 11:35:24 2016 |
steve | Update | safety | BS visitor's viewport is protected with lexan |
Quote: | The four horizontal viewports of arms are protected
by 3/8" thick, 8.5" OD Lexan disk of MR10 Polycarbonate.
ITMX, ETMX, ITMY and ETMY ccd cameras are not focused now. |
BS visitor's viewport glass is now covered with Lexan MR10
Note:
this Lexan cover is in vertical orientation so becomes lose when the black anodized cover is removed.
It needs to be held in place while it's housing is taken off. |
12053
|
Tue Mar 29 03:16:21 2016 |
ericq | Update | LSC | DRFPMI Locked Once More |
[ericq, Gautam]
Three RF-only locks longer than a minute tonight, out of 5 total attempts.
Last week, I determined that the beam spot on the RF POP PD is too large. This still needs to be fixed. I updated the ASS model to use REFLDC as a PRCL dither error signal; it works.
There seems to be some excess angular motion of ETMY tonight. This is evident in the oplev spectra (as compared to ETMX), and the GTRY camera, and even the retroreflected beam from a misalgined ETMY on the ITMY face when the PRC is carrier locked.
Gautam and I mostly focused on setting up the CAL-DARM_CINV block to produce this (mostly) calibrated spectrum starting from GPS 1143274087. [Darm on unwhitened AS55, DRMI on 3F, one CARM boost]

Here are the control and error signal spectra:

[DTT files attached]
Note to self: archive some of this data |
Attachment 1: 2016-03-29_calibdarm.pdf
|
|
Attachment 2: 2016-03-29_DRFPMI_errctrl.pdf
|
|
Attachment 3: DRFPMI_DTT.zip
|
12052
|
Mon Mar 28 22:16:44 2016 |
Koji | Update | General | New WiFi router |
I configured three more mini wifi extender. They are ready to use.
We should add these to the host table (I forgot where it is)
192.168.113.233 NETGEAR_EX3700_1
192.168.113.234 NETGEAR_EX3700_2
192.168.113.235 NETGEAR_EX3700_3
192.168.113.236 NETGEAR_EX3700_4
|
12051
|
Mon Mar 28 10:43:18 2016 |
Steve | Update | endtable upgrade | ETMX 4'x2' optical table pictures |
|
Attachment 1: ETMX4x2Layout.jpg
|
|
Attachment 2: ETMX_4x2.JPG
|
|
Attachment 3: ETMXopt_4x2.JPG
|
|
12050
|
Mon Mar 28 08:30:09 2016 |
Steve | Update | SUS | PRM damping restored |
Recent EQ 4.8 mag San Felipe, Mexico trips PRM sus damping.
PRM damping restored. PMC locked. |
Attachment 1: shaking.png
|
|
12049
|
Sat Mar 26 18:28:24 2016 |
Koji | Update | elog | elogd flakiness |
Elogd have been restarted several times today because it died everytime I submit something.
Here is the copy of the log.
GET /OMC_Lab/255?cmd=loc&value=Submit HTTP/1.1
Returned 6 bytes
GET /40m/elog.rdf HTTP/1.1
Returned 17109 bytes
TCP connection #1 on socket 5 closed
POST /OMC_Lab/ HTTP/1.1
Returned 20 bytes
GET /OMC_Lab/255 HTTP/1.1
Returned 53721 bytes
GET /ckeditor/skins/moono/images/arrow.png HTTP/1.1
Returned 489 bytes
POST /OMC_Lab/ HTTP/1.1
*** buffer overflow detected ***: /export/home/elog/elog/elogd terminated
======= Backtrace: =========
/lib/x86_64-linux-gnu/libc.so.6(__fortify_fail+0x37)[0x7f1435639e57]
/lib/x86_64-linux-gnu/libc.so.6(+0x108d50)[0x7f1435638d50]
/lib/x86_64-linux-gnu/libc.so.6(+0x1081b9)[0x7f14356381b9]
/lib/x86_64-linux-gnu/libc.so.6(_IO_default_xsputn+0xdd)[0x7f14355ab0cd]
/lib/x86_64-linux-gnu/libc.so.6(_IO_vfprintf+0x25a8)[0x7f143557ac18]
/lib/x86_64-linux-gnu/libc.so.6(__vsprintf_chk+0x94)[0x7f1435638254]
/lib/x86_64-linux-gnu/libc.so.6(__sprintf_chk+0x7d)[0x7f143563819d]
/export/home/elog/elog/elogd[0x426405]
/export/home/elog/elog/elogd[0x473b7f]
/export/home/elog/elog/elogd[0x4abfb2]
/export/home/elog/elog/elogd[0x4ad7fb]
/export/home/elog/elog/elogd[0x4b0af5]
/export/home/elog/elog/elogd[0x4b1eb9]
/export/home/elog/elog/elogd[0x403568]
/lib/x86_64-linux-gnu/libc.so.6(__libc_start_main+0xed)[0x7f143555176d]
/export/home/elog/elog/elogd[0x404299]
======= Memory map: ========
00400000-004e6000 r-xp 00000000 fc:00 10361276 /export/home/elog/elog-3.0.d/elogd
006e5000-006e6000 r--p 000e5000 fc:00 10361276 /export/home/elog/elog-3.0.d/elogd
006e6000-007c6000 rw-p 000e6000 fc:00 10361276 /export/home/elog/elog-3.0.d/elogd
007c6000-0173d000 rw-p 00000000 00:00 0
0214d000-02656000 rw-p 00000000 00:00 0 [heap]
7f14342f8000-7f143430d000 r-xp 00000000 fc:00 2883628 /lib/x86_64-linux-gnu/libgcc_s.so.1
7f143430d000-7f143450c000 ---p 00015000 fc:00 2883628 /lib/x86_64-linux-gnu/libgcc_s.so.1
7f143450c000-7f143450d000 r--p 00014000 fc:00 2883628 /lib/x86_64-linux-gnu/libgcc_s.so.1
7f143450d000-7f143450e000 rw-p 00015000 fc:00 2883628 /lib/x86_64-linux-gnu/libgcc_s.so.1
7f143450e000-7f14348cd000 rw-p 00000000 00:00 0
7f1434a34000-7f1434d39000 r--p 00000000 fc:00 530477 /usr/lib/locale/locale-archive
7f1434d39000-7f1434d4f000 r-xp 00000000 fc:00 655527 /usr/local/lib/libz.so.1.2.8
7f1434d4f000-7f1434f4e000 ---p 00016000 fc:00 655527 /usr/local/lib/libz.so.1.2.8
7f1434f4e000-7f1434f4f000 r--p 00015000 fc:00 655527 /usr/local/lib/libz.so.1.2.8
7f1434f4f000-7f1434f50000 rw-p 00016000 fc:00 655527 /usr/local/lib/libz.so.1.2.8
7f1434f50000-7f1434f52000 r-xp 00000000 fc:00 2883655 /lib/x86_64-linux-gnu/libdl-2.15.so
7f1434f52000-7f1435152000 ---p 00002000 fc:00 2883655 /lib/x86_64-linux-gnu/libdl-2.15.so
7f1435152000-7f1435153000 r--p 00002000 fc:00 2883655 /lib/x86_64-linux-gnu/libdl-2.15.so
7f1435153000-7f1435154000 rw-p 00003000 fc:00 2883655 /lib/x86_64-linux-gnu/libdl-2.15.so
7f1435154000-7f1435307000 r-xp 00000000 fc:00 2883609 /lib/x86_64-linux-gnu/libcrypto.so.1.0.0
7f1435307000-7f1435506000 ---p 001b3000 fc:00 2883609 /lib/x86_64-linux-gnu/libcrypto.so.1.0.0
7f1435506000-7f1435521000 r--p 001b2000 fc:00 2883609 /lib/x86_64-linux-gnu/libcrypto.so.1.0.0
7f1435521000-7f143552c000 rw-p 001cd000 fc:00 2883609 /lib/x86_64-linux-gnu/libcrypto.so.1.0.0
7f143552c000-7f1435530000 rw-p 00000000 00:00 0
7f1435530000-7f14356e4000 r-xp 00000000 fc:00 2884139 /lib/x86_64-linux-gnu/libc-2.15.so
7f14356e4000-7f14358e3000 ---p 001b4000 fc:00 2884139 /lib/x86_64-linux-gnu/libc-2.15.so
7f14358e3000-7f14358e7000 r--p 001b3000 fc:00 2884139 /lib/x86_64-linux-gnu/libc-2.15.so
7f14358e7000-7f14358e9000 rw-p 001b7000 fc:00 2884139 /lib/x86_64-linux-gnu/libc-2.15.so
7f14358e9000-7f14358ee000 rw-p 00000000 00:00 0
7f14358ee000-7f1435943000 r-xp 00000000 fc:00 2884155 /lib/x86_64-linux-gnu/libssl.so.1.0.0
7f1435943000-7f1435b42000 ---p 00055000 fc:00 2884155 /lib/x86_64-linux-gnu/libssl.so.1.0.0
7f1435b42000-7f1435b45000 r--p 00054000 fc:00 2884155 /lib/x86_64-linux-gnu/libssl.so.1.0.0
7f1435b45000-7f1435b4c000 rw-p 00057000 fc:00 2884155 /lib/x86_64-linux-gnu/libssl.so.1.0.0
7f1435b4c000-7f1435b6e000 r-xp 00000000 fc:00 2884145 /lib/x86_64-linux-gnu/ld-2.15.so
7f1435d57000-7f1435d5c000 rw-p 00000000 00:00 0
7f1435d6a000-7f1435d6e000 rw-p 00000000 00:00 0
7f1435d6e000-7f1435d6f000 r--p 00022000 fc:00 2884145 /lib/x86_64-linux-gnu/ld-2.15.so
7f1435d6f000-7f1435d71000 rw-p 00023000 fc:00 2884145 /lib/x86_64-linux-gnu/ld-2.15.so
7ffd85795000-7ffd85997000 rw-p 00000000 00:00 0 [stack]
7ffd859b2000-7ffd859b4000 r-xp 00000000 00:00 0 [vdso]
ffffffffff600000-ffffffffff601000 r-xp 00000000 00:00 0 [vsyscall]
er_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "cache_disable"
Received unknown cookie "NO_CACHE"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id" |
12048
|
Fri Mar 25 23:54:04 2016 |
rana | Update | PEM | Guralp Seismometers |
Something seems not right. The Guralp response should be flat in velocity from 0.05-30 Hz. Why is there any feature at 1 Hz? Saturation of some kind? |
12047
|
Fri Mar 25 19:17:28 2016 |
Nikhil | Update | PEM | Guralp Seismometers |
Calibration of Guralp Seismometers
Objective
- Estimate transfer functions of Guralp A ( near ETMX) and Guralp B ( near ETMY)
- Calibrate the instruments by estimating Velocity Sensitity Parameter
- Convert previously measured Voltage Spectrum to Velocity Spectrum
Instruments Used
- Guralp CMG-40 T Seimometers : Guralp A (Serial Number: T4Q17)
- Guralp CMG-40 T Seimometers : Guralp B (Serial Number: T4157)
- Guralp Handheld Control Unit (HCU)
- FFT Spectrum Analyzer: Model SR785: 2 Channel Dynamic Signal Analyzer
- Oscilloscope: TDS 3014B
- Function Generator: DS 345
Procedure & Results
Sinusoidal current of known frequency and amplitude was injected to the Seismometer calibration coil using signal generator and handheld control unit & corresponding Magnitude and Phase response were recorded. For Guralp B, system response was also estimated with a FFT Spectrum Analyzer.

Frequnecy Range: 0.1 Hz to 45 Hz.
Equivalent Input Velocity was derived from the Input Voltage measurements using the relation: v = V/ (2*pi*f*R*K) , V is the peak to peak Calibration Signal voltage, f is the calibration signal frequency, R is the calibration resistor and K is the feedback coil constant. [See Appendix for R & K values]
Velocity Sensitity at the required frequency is obtained by dividing the Output Response Voltage by the Equivalent Input Velocity.

The obtained Velocity Sensitivity is used to convert the recorded Volatge PSD to Velocity PSD as shown below. The obtained results are compared to gloabl high noise model [NHNM] and USGS New Low Noise Model [NLNM,Peterson 1993] which gives the lowest observed vertical seismic noise levels across the seismic frequency band. Plot legend NLNM shows both the high & low levels.
Guralp A [X Arm] Low Velocity Output

Guralp B [Y Arm] Low Velocity Output

DTT Power Spectrum
Both the Seismometers were connected to the 40 M Control and Data Acquisition System (CDS) and Power Spectrum was estimated for the Vertical, North/South & East/West Channels using Diagnostic Test Tool (DTT) software.

Comments
- The transfer function from Guralp A [ETMX] looks similar to that of Guralp B [ETMY] in both magnitude and phase but with a lower gain.
- Velocity Sensitivity of Guralp A is comparable to the value provided in the Calibration Data Sheet [~ 400] for all the channels [Vertical, North/South, East/West] after 1 Hz. For Guralp B, Velocity Sensitivity is a factor of 2.5 higher [all channels] than the specification [~ 400] after 1 HZ.Below 1 Hz Sensitivity drops down for both sensors. I am not ruling out a missing common factor in the calculation, but anyway, test shows that Guralp B has ~2.5 times better Velocity Sensitivity than Guralp A.
- The Calibrated Seismic Velocity Spectrum for Guralp B is within the Globally Observed High and Low Noise Seismic Spectrum while Guralp A's Spectrum is more noisier above 1 Hz [Anthropogenic Activity normally contributes the most in 1 Hz to 10 Hz frequency band].
- Concurrently acquired Power Spectrum using DTT [Diagnostic Test Tools] shows that Guralp A Spectrum behaves rather strangely. The system response seems to be completely different from the one we obtained locally using signal generator. While Guralp B functionality seems normal. One reason for this erratic beahvior might be faulty cables used for data acquisition from Guralp A. This needs to be verified.
Appendix
CMG-40T Guralp A Calibration Sheet
|
Velocity Output: V/m/s (Differential) |
Mass Position Output (Acceleration Output) : V/m/s^2 |
Feedback Coil Constant : Amp/m/s^2 |
Vertical |
2 x 400 |
19 |
0.00397 |
North/South |
2 x 398 |
23 |
0.00486 |
East/West |
2 x 401 |
23 |
0.00479 |
Calibration Resistor: 51000
CMG-40T Guralp B Calibration Sheet
|
Velocity Output: V/m/s (Differential) |
Mass Position Output (Acceleration Output) : V/m/s^2 |
Feedback Coil Constant : Amp/m/s^2 |
Vertical |
2 x 401 |
19 |
0.00408 |
North/South |
2 x 400 |
20 |
0.00421 |
East/West |
2 x 404 |
22 |
0.00466
|
Calibration Resistor: 51000 |
Attachment 3: VelSens_XArm_Guralp_A.png
|
|
Attachment 4: VelSens_XArm_Guralp_A.png
|
|
Attachment 5: VelSens_YArm_Guralp_B.png
|
|
Attachment 8: Vel_PSD_XArm_GurB_E.png
|
|
Attachment 9: Vel_PSD_XArm_GurB_N.png
|
|
Attachment 16: Guralp_PowerSpectrum.pdf
|
|
12046
|
Thu Mar 24 08:20:52 2016 |
Steve | Update | PEM | Guralp-A calibration sheet |
All Guralp instruments and digitisers are provided with calibration documentation. Should you require a copy of calibration information for any product, email caldoc@guralp.com with the serial number of the product in the subject field and calibration information will be sent to you through email.
See data in the 40m wiki
|
12045
|
Thu Mar 24 07:56:09 2016 |
Steve | Update | Calibration-Repair | NO Noise Eater for 1W Innolight |
1W Innolight is NOT getting Noise Eater as it was decided yesterday at the 40m meeting. Corrected 3-25-2016
Repair quote with adding noise eater is in 40m wiki
Quote: |
Quote: |
Quote: |
After adjusting the alignment of the two beams onto the PD, I managed to recover a stronger beatnote of ~ -10dBm. I managed to take some measurements with the PLL locked, and will put up a more detailed post later in the evening. I turned the IMC autolocker off, turned the 11MHz Marconi output off, and closed the PSL shutter for the duration of my work, but have reverted these to their nominal state now. The are a few extra cables running from the PSL table to the area near the IOO rack where I was doing the measurements from, I've left these as is for now in case I need to take some more data later in the evening...I
|
Innolight 1W 1064nm, sn 1634 was purchased in 9-18-2006 at CIT. It came to the 40m around 2010
It's diodes should be replaced, based on it's age and performance.
RIN and noise eater bad. I will get a quote on this job.
The Innolight Manual frequency noise plot is the same as Lightwave' elog 11956
|
Diagnoses from Glasglow:
“So far we have analyzed the laser. The pump diode is degraded. Next we would replace it with a new diode. We would realign the diode output beam into the laser crystal. We check all the relevant laser parameters over the whole tuning range. Parameters include single direction operation of the ring resonator, single frequency operation, beam profile and others. If one of them is out of spec, then we would take actions accordingly. We would also monitor the output power stability over one night. Then we repackage and ship the laser.”
|
|
12044
|
Wed Mar 23 15:23:12 2016 |
Steve | Summary | PEM | Guralps as connected |
We have one calibration sheet of GUR- B, from 26 June 2008, model CMG-T40-0008, sn T4157 at ETMY east, interface box input 1
I'm looking for calibration paper of GUR- A, model CMG-T40-0053, sn T4Q17 at ETMX south, interface box input 2
Quote: |
I measured the guralp raw outputs and the TFs using the handheld unit and an FFT analyzer.
[Setup]
The handheld unit was connected to each guralp with the same cable which is confirmed t be functional with the Yend Guralp.
The signal for Z, N, and E directions are obtained from the banana connectors on the handheld unit. Each direction has mass, low gain velocity, and high gain velocity output. The PSDs of the signals were measured with an FFT analyzer. The transfer function from the mass signal to the low/high gain signals were also measured for each direction.
The adjustment screw for the E output of the Xend does not work. I had to tilt the Xend Guralp using the leg screws to bring the E signal to zero.
[Result]
Attachment 1: Raw voltage PSD for all outputs
Attachment 2: Comparison of the low gain vel outputs
- All of the mass output show similar PSDs.
- Low gain velocity outputs shows somewhat similar levels. I still need to check if the output is really the ground velocity or not.
- High gain velocity outputs are either not high gain, broken, or not implemented.
- We need to calibrate the low gain output using signal injection, huddle test, or something else.
Attachment 3: TFs between each mass output and the low or high gain outputs
- TFs between the mass signal and the low vel signals show the similar transfer functions between the channels.
- The high gain outputs show low or no transfer function with regard to the mass signals.
|
|
Attachment 1: GUR_A.jpg
|
|
Attachment 2: GUR_B.jpg
|
|
Attachment 3: GUR_Interface_Box.jpg
|
|
12043
|
Wed Mar 23 11:55:47 2016 |
Steve | Update | General | Smart UPS 2200 Battery Replaced |
Batteries replaced in control room UPS after 3 years from replaceUPSbattery.com
|
12042
|
Tue Mar 22 21:30:15 2016 |
Koji | Update | IOO | PMCIMC aligned, WFS offset adjusted |
The alignment of the PMC adjusted on the PSL table: Trans 0.737->0.749
The alignment of the IMC adjusrted on the sliders: Trans 14300->15300
WFS offset has been reset by /opt/rtcds/caltech/c1/scripts/MC/WFS/WFSoffsets |
Attachment 1: 08.png
|
|
12041
|
Tue Mar 22 14:12:18 2016 |
Steve | Update | IOO | Laser is turned back on |
The 2W Innolight was turned on.
|
Attachment 1: off-onAgain.png
|
|
12040
|
Mon Mar 21 14:29:32 2016 |
Steve | Update | Calibration-Repair | 1W Innolight laser repair diagnoses |
Quote: |
Quote: |
After adjusting the alignment of the two beams onto the PD, I managed to recover a stronger beatnote of ~ -10dBm. I managed to take some measurements with the PLL locked, and will put up a more detailed post later in the evening. I turned the IMC autolocker off, turned the 11MHz Marconi output off, and closed the PSL shutter for the duration of my work, but have reverted these to their nominal state now. The are a few extra cables running from the PSL table to the area near the IOO rack where I was doing the measurements from, I've left these as is for now in case I need to take some more data later in the evening...I
|
Innolight 1W 1064nm, sn 1634 was purchased in 9-18-2006 at CIT. It came to the 40m around 2010
It's diodes should be replaced, based on it's age and performance.
RIN and noise eater bad. I will get a quote on this job.
The Innolight Manual frequency noise plot is the same as Lightwave' elog 11956
|
Diagnoses from Glasglow:
“So far we have analyzed the laser. The pump diode is degraded. Next we would replace it with a new diode. We would realign the diode output beam into the laser crystal. We check all the relevant laser parameters over the whole tuning range. Parameters include single direction operation of the ring resonator, single frequency operation, beam profile and others. If one of them is out of spec, then we would take actions accordingly. We would also monitor the output power stability over one night. Then we repackage and ship the laser.” |
12039
|
Mon Mar 21 10:07:45 2016 |
Steve | Update | SUS | ETMX enclosure update |
The enclosure top piece in the middle is still in the machine shop.
The carpenter who helps in the built just left for one week vacation.
The unit will be ready on April 1
I'd prefer doing the installation with the enclosure on the new table.
It's the only way to minimize the resonances of the enclosure with shimming. |
Attachment 1: ETMXenclosure.jpg
|
|
12038
|
Fri Mar 18 11:04:56 2016 |
Steve | Update | SUS | inverted pendulum |
Kate Dooley picked up this item today. |
Attachment 1: invPeng.jpg
|
|
12037
|
Wed Mar 16 16:02:40 2016 |
Koji | Update | SUS | wire standoff test cut |
It looks almost OK, but we need a bit sharper picture for both the groove and thw wire. |
Attachment 1: ruby.png
|
|
12036
|
Wed Mar 16 15:36:03 2016 |
Steve | Update | SUS | wire standoff test cut |
Ruby wire standoff 1 mm od. with V-groove test cut. SOS sus wire 0.0017" od. is in the background.
|
Attachment 1: 0182_-_20160316_135935.png
|
|
12035
|
Tue Mar 15 10:31:58 2016 |
Steve | Update | IOO | Laser is turned back on |
It's may be the janitor's doing.
I noticed that the HEPA filers were off. They are turned on at 20%
|
Attachment 1: 2WlaserOff-On.png
|
|
12033
|
Mon Mar 14 22:42:23 2016 |
gautam | Update | endtable upgrade | Inventory check |
Quote: |
Steve should be able to get another copy of the EY doubler mount made up if we really don't have another one sitting in the Manasa end table box which Koji mentioned.
|
I located the second doubler mount, it was sitting inside a cabinet along the Y-arm. So this will not have to be machined. The doubling oven mount is black in colour.
So as things stand now, the only thing that needs to be machined is a non-green mount for the IR faraday (IO-5-1064-HP) - is it possible to just coat the existing mount with a different color? I've got a drawing for this part ready, but it seems unnecessary to machine the whole thing from scratch when only the color is an issue. Steve was talking about dipping this in some sort of solution and taking the green off. But if this isn't possible, I'll send Steve the drawings tomorrow so that he can place the order with the machine shop...
I will work on the mode-matching calculations over the next couple of days to make sure we have all the mirrors and lenses we need.
|
12032
|
Sat Mar 12 22:23:37 2016 |
rana | Summary | IOO | PMC relocked |
Found it locked on TEM01 mode.
Sweets in the fridge for non-PhD holders, courtesy of the highest levels of Caltech. |
12031
|
Fri Mar 11 16:52:53 2016 |
Steve | Update | PEM | leaky roof |
Johannes found dripping water at the vac rack. It is safe. It is not catching anything. Actual precipitation was only 0.62" |
Attachment 1: leakyRoof.jpg
|
|
Attachment 2: leakyRoofA.jpg
|
|
12030
|
Thu Mar 10 16:32:45 2016 |
rana | Update | endtable upgrade | Inventory check |
Its not a good idea to use green mounts with green lasers. Steve should be able to get another copy of the EY doubler mount made up if we really don't have another one sitting in the Manasa end table box which Koji mentioned. |
12029
|
Thu Mar 10 16:29:32 2016 |
gautam | Update | endtable upgrade | Inventory check |
I did a quick sweep of the lab to find out what hardware has already been acquired for the X-end table upgrade. The attached PDF is an inventory check in the spirit of this elog.
Some things we have to decide:
- Are we okay with using the old green coloured faraday mount for the IR faraday? I have in hand a piece identical to the one used at the Y-end for the green faraday, that is red in colour, so I guess we can switch this out.
- The way in which the doubling oven is currently mounted at the X-end is using some posts cobbled together. The Y-end looks to have a custom mount machined for it (see Attachment #2). Do we want to go ahead and get something like this done?
- I suppose it is okay to reuse all the old optics (mirrors, lenses, harmonic separators) and PDs? It may be that we need to order som extra mirrors/lenses/posts (this will become clear once I do the layout)
I have not gotten around to planning the layout or doing drawings. I will try and first work through a mode-matching solution to make sure we have all the required lenses. It may be that we need some 1" or 2" mirrors as well. The beam from the lightwave NPRO is quite elliptical, but we have a number of cylindrical lenses in hand already if we decide we want to use these, so I guess we don't have to worry about this...
This is quite a preliminary list, and I will add/update over the coming days as I do more detailed planning, but have I missed out anything obvious? |
Attachment 1: Inventory_check.pdf
|
|
Attachment 2: Doubler_comparison.pdf
|
|
12028
|
Thu Mar 10 03:03:11 2016 |
ericq | Update | LSC | DRFPMI Power stable, but no RF handoff |
[ericq, Gautam]
We worked on getting the DRFPMI back up and running, hoping the ALS performance was good enough.
We did succeed in bringing in enough of the AO path to stabilize arm powers > 100, but failed at the full RF DARM handoff.
REFL165 angle was adjusted to -86 to minimize PRCL in the Q signal.
The AS110 signals are mysteriously huger than they used to be. Whitening gain reduced to 15dB from 27dB. Old trigger thresholds are still fine.
The new AUX X laser has a different sign for the temperature-> frequency coupling, so our usual convention of "beatnote goes up when temp slider goes up" meant the ALSX input matrix elements had to change sign.
We think the POPDC PD (which I think is the POP2F PD) may be miscentered, since in PRMI configuration, its maximum does not coincide with the REFLDC minimum, and leaves a sizeable TEM10 lobe on the REFL camera. This was a pain. |
12027
|
Tue Mar 8 18:22:20 2016 |
rana | Update | Green Locking | Laser swap - some improvement |
Why is the transmission of X green so low? Perhaps you can phase lock the IR and then scan the X frequency, using the X arm as the analyzer. i.e. put a slow ramp into MC2 to pull the PSL frquency and thus the green frequency. You can record a movie of the scan using the framegrabber and record the green transmission peaks to see how big the mode match is exactly (which modes are so big) |
12026
|
Mon Mar 7 23:51:36 2016 |
gautam | Update | Green Locking | Laser swap - some improvement |
Quote: |
Next steps in recovering ALS and trying to lock again
- Having set the PDH modulation frequency to 256.62kHz, I took the spectrum of ALS noise using the IR beat (i.e. by piping the IR beat signal through the electronics the green beats usually go through - 6dB and 10dB attenuators were placed immediately after the beat PDs for the X and Y arms respectively, to make the signal levels compatible with the electronics), Attachment #5 unfortunately suggests that the noise performance is still poor, and I suspect the situation will be similar using the green beat (though I have not measured this yet).
- The modulation depth could be sub-optimal for the X-end PDH, I have to measure this and check that it is at an acceptable level. This will also tell me if I need to change the sum+HPF pomona box used to send the PDH control signal + piezo dither signal to the laser PZT. In order to do this, I need to know what the input impedance to the FAST control BNC is - the manual isn't very helpful, it just says the piezo has a capacitance less than 10,000pF. I suppose I will have to actually measure this.
- PDH loop OLTFs have to be re-measured for both ends to check that the servo gain's are appropriately placed.
- We know that the mode-matching into the arm for the X end is poor (I have yet to quantify this) - I suspect that the beam ellipticity is the main culprit. However, the DC transmitted power levels at the PSL table are comparable to (even slightly better than) the Y arm numbers, and so this cannot be the sole reason why the X-arm ALS noise is so much worse... I will continue my investigations next week...
|
Attachment #1
Since I could not determine how many volts at the LO input of the pomona box input corresponds to how many volts at the laser PZT, I measured the transfer function between these points using the Agilent network analyzer. The measured TF suggests that for a function generator output of 2Vpp, we get approximately 75mrad of phase modulation, which compares reasonably well with the value of 120mrad reported here. I did not attempt to further increase the LO output signal to push this number closer to 120mrad, as with 2Vpp from the function generator we get +7dBm at the mixer, which is what it wants - so I wanted to avoid any attenuators etc...
Attachments #2 and #3
After ensuring that we have appreciable phase modulation, I set out to measure the PDH OLTFs and adjust the gain on the uPDH boxes accordingly. The X end gain is at 6.0, and the Y end gain is at 4.0. Before measuring the Y-end OLTF, I adjusted the steering mirrors to increase GTRY to ~0.45. GTRX remains a paltry 0.05... But the UGFs seem satisfactory..
Attachment #4
Finally, I took the ALS noise spectrum for the green beats. The beat note amplitudes on the network analyzer in the control room are still puny compared to what we had, -40dBm for Y and -45dBm for X. But the phase tracker Q values are ~1000 and ~3000 for X and Y respectively, which are pretty close to what these were if memory serves me right. There may still be some room for optimization of the PDH loop gains etc, and we could perhaps look at lowering the gain of the REFL PD at the X end? I also have yet to do the sweep for the 3 temperatures at which we can find a beatnote and park at the middle one...
These spectra suggest we could even possibly try locking? We are approximately a factor of 3 above the reference for X and on par with the reference for Y....
Unrelated to this work: I also realinged the PMC, PMC transmission is now 0.730V up from ~0.65V. |
Attachment 1: PomonaTF.pdf
|
|
Attachment 2: XPDH.pdf
|
|
Attachment 3: YPDH.pdf
|
|
Attachment 4: greenbeat_20160307.pdf
|
|
12025
|
Mon Mar 7 20:40:02 2016 |
ericq | Update | CDS | FB down again |
We went and looked at the monitor plugged into FB. All kinds of messages were being spammed to the screen (maybe RAM errors), and nothing could be done to interrupt. Sadly, a hard reboot of FB was neccesary.
Video of error messages: https://youtu.be/7rea_kokhPY
After the reboot, it just took a couple of model restarts to get the CDS screen happy. |
12024
|
Sun Mar 6 15:24:05 2016 |
gautam | Update | CDS | FB down again |
I came in to check the status of the nitrogen and noticed that the striptool panels in the control room were all blank.
- PMC was unlocked but I was able to relock it using the usual procedure
- FB seems to be down: I was unable to ssh into it (or any of the FEs for that matter). I checked the lights on the RAID array, they are all green. I am holding off on doing a hard reboot of FB in case there is some other debugging that can be done first
- None of the watchdogs were tripped, but judging by the green spots on the mirrors, all of them are moving quite a bit. I've shutdown the watchdogs on all the optics except the MC mirrors, but the ITMs and ETMs still seem to be moving quite a bit.
I am leaving things in this state for now. It is unclear why this should have happened, it doesn't seem like there was a power glitch? |
Attachment 1: 58.png
|
|
12023
|
Sat Mar 5 23:31:01 2016 |
gautam | Update | Green Locking | Laser swap - some updates |
I've been a little behind on my elogs so here is an update of the end laser situation.
IR beat for X-end recovered
- The issue was optimizing the alignment into the fiber at the end table.
- Using Fluke fiber illuminator helped in aligning IR pickoff into mount. Useful note: there is an unused fiber running between the X-end and the PSL table, by connecting these at the PSL table, I was able to monitor the coupled power while remaining at the X-end.
- Another major issue was that one of the steering mirrors (marked "Y1" in Attachment #1) was mounted with AR coated side facing the beam. This was fixed by simply rotating the post, the mirror was not removed from its mount. I can only assume that this mirror is in this kind of mount because of space constraints.
- The fiber has a collimating telescope attached to the end of it. In principle, this gives us more angular acceptance while coupling the beam into the fiber, but as I found out, the acceptance is still tiny (I don't have a number to quantify it). Furthermore, the Fluke visual fault locator revealed that the lens in the collimating telescope is not set up great - when re-doing the X end table, we should fix this situation so as to have a fairly large collimated beam coming out of the fiber when illuminated from the other end, this would make the mode matching much easier.
- Bottom line: we have ~1.2 mW of IR light incident on the coupler at the end table, and ~400uW of IR power at the PSL table => coupling efficiency is ~30%, not stellar, but sufficient for now I guess. After the various splitters etc, there is about 160uW of EX IR light and ~300uW of PSL IR light incident on the beat PD, and the beat amplitude is about -9dBm.
AM/PM characterization of newly installed Lightwave
- Having recovered the IR beat, I set out to do the PM characterization for the end laser.
- Attachment #2 shows the electrical setup. The IR beat was piped to the X-end via an existing long cable that runs between the vertex and the endtable. Not shown in the diagram, but I used a 20dB coupler to keep track of the beat frequency on the HP spectrum analyzer while doing this measurement.
- I restricted myself to the range between 100kHz and 500kHz to do the scan, because it takes quite a while to do the scan with fine resolution (IF bandwidth = 10Hz).
- To calibrate the magnitude response to rad/V, I divided the output of the network analyzer (converting dB to absolute magnitude first) by the amplitude of the signal seen on the monitoring oscilloscope while the PLL is unlocked. This number was 96mV/rad.
- To confirm that the error signal spectrum is indeed a good approximation of the "plant" transfer function (i.e that 100kHz >> UGF of loop transfer function of the PLL), I measured the loop TF of the PLL - Attachment #3 suggests a UGF of ~ 16kHz, which means the assumption is reasonable.
- Excitation amplitude was -25dBm (which gave reasonable SNR), and 3 averages were taken.
- The AM measurement was done using the same procedure as detailed here - the DC block was used. The DC level of the PD output was 2.72 V. The excitation amplitude was 0dBm.
- Attachment #4 shows the AM response, PM response and PM/AM ratio
- The peak in the PM/AM ratio at 256620 Hz is compelling because it is not too sharp (and so we can be reasonably confident we are at a good operating point) and the PM response of 23.83 rad/V is also acceptable.
- As a consistency check, the PM response of ~30rad/V at 100kHz => PZT actuator gain is ~3MHz/V, which is in the region we expect it to be...
Next steps in recovering ALS and trying to lock again
- Having set the PDH modulation frequency to 256.62kHz, I took the spectrum of ALS noise using the IR beat (i.e. by piping the IR beat signal through the electronics the green beats usually go through - 6dB and 10dB attenuators were placed immediately after the beat PDs for the X and Y arms respectively, to make the signal levels compatible with the electronics), Attachment #5 unfortunately suggests that the noise performance is still poor, and I suspect the situation will be similar using the green beat (though I have not measured this yet).
- The modulation depth could be sub-optimal for the X-end PDH, I have to measure this and check that it is at an acceptable level. This will also tell me if I need to change the sum+HPF pomona box used to send the PDH control signal + piezo dither signal to the laser PZT. In order to do this, I need to know what the input impedance to the FAST control BNC is - the manual isn't very helpful, it just says the piezo has a capacitance less than 10,000pF. I suppose I will have to actually measure this.
- PDH loop OLTFs have to be re-measured for both ends to check that the servo gain's are appropriately placed.
- We know that the mode-matching into the arm for the X end is poor (I have yet to quantify this) - I suspect that the beam ellipticity is the main culprit. However, the DC transmitted power levels at the PSL table are comparable to (even slightly better than) the Y arm numbers, and so this cannot be the sole reason why the X-arm ALS noise is so much worse... I will continue my investigations next week...
|
Attachment 1: AUXxTelescope.png.png
|
|
Attachment 2: PM_setup.pdf
|
|
Attachment 3: PLLolg.pdf
|
|
Attachment 4: AMPM20160303.pdf
|
|
Attachment 5: IRbeat_20160304.pdf
|
|